Grade 11Trigonometry

What`s the value of sin217 in fractions???Please tell the answer

Nischal C N
9 Years agoGrade 11

1 Answer

Saurabh Koranglekar
6 Years ago
Dear student

It is sin (180 + 37) = -sin(37) = -3/5

Regards